OpenStudy (anonymous):

8 th grade math... The length of a slide of an equilateral triangle is represented by x+5. Represent the perimeter of the triangle

OpenStudy (saifoo.khan):

(x+5) + (x+5) + (x+5)

OpenStudy (saifoo.khan):

Solve^

OpenStudy (anonymous):

3x+15

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

can i?

OpenStudy (saifoo.khan):

As it is an equilateral triangle, all the sides will be the same.

OpenStudy (saifoo.khan):

As some is x + 5, so other two will be the same.

OpenStudy (anonymous):

oh, so X=5?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

Nops, it can't be that. this is b/c the question gives us an expression, not an equation.

OpenStudy (anonymous):

one side = x+5 perimeter = sum of all side =sum of three sides = x+5 + x+5 +x + 5

OpenStudy (anonymous):

=3x+15

OpenStudy (anonymous):

ok